Find the solution to y + 2y – 8y =0.
Answer #1

y - 27-8y=0 -> (p?.+20- 8 )y=0 whese Da auxiliary equation in d an The m? +2m 8 - 0 > m² + 4m- 2m - 8=0 cm (m+4)=21m+4)=0 ->